Calculate the present value of each of the following cash flows using a discounting rate of 12%
a. $6000 cash inflow at the end of each of the next 6 years.
b. $6000 cash inflow at the beginning of each of the next 6 years.
1
Expert's answer
2020-09-21T15:51:48-0400

present value of inflow = cash inflow * present value discounted factor

a. At the end of each year

1st year ="\dfrac{6000}{1.12}" =5357.1428

2ndyear ="\dfrac{6000}{1.12^2}" =4783.163

3rd year ="\dfrac{6000}{1.12^3}" =4270.681

4thyear ="\dfrac{6000}{1.12^4}" =3813.108

5th year ="\dfrac{6000}{1.12^5}" =3404.561

6thyear ="\dfrac{6000}{1.12^6}" =3039.7867

total ="\dfrac{6000}{1.12}+\dfrac{6000}{1.12^2}+\dfrac{6000}{1.12^3}+\dfrac{6000}{1.12^4}+\dfrac{6000}{1.12^5}+\dfrac{6000}{1.12^6}"

=$24668.44



b. At the beginning of each year there will be addition of $6000 at the start

1st year =6000

2ndyear ="\dfrac{6000}{1.12^2}" =4783.163

3rd year ="\dfrac{6000}{1.12^3}" =4270.681

4thyear ="\dfrac{6000}{1.12^4}" =3813.108

5th year ="\dfrac{6000}{1.12^5}" =3404.561

6thyear ="\dfrac{6000}{1.12^6}" =3039.7867

"\therefore" PV ="6000+\dfrac{6000}{1.12^2}+\dfrac{6000}{1.12^3}+\dfrac{6000}{1.12^4}+\dfrac{6000}{1.12^5}+\dfrac{6000}{1.12^6}"

=$25311.30